Find the equation of the line through the origin with slope -6 Answer: 6x+y=0
1
Expert's answer
2020-03-03T14:00:39-0500

The equation of the line with the slope has the form

"y=kx+b"

According to the problem, "k=\tan\alpha=-6,"

then "y=-6x+b" .

The line goes through the origin  "O(0,0)" , hence

substitute coordinates of the origin into the equation of the line:

"0=-6\cdot0+b,\\
b=0."

The equation of the line is

"y=-6x,\\
6x+y=0."